OpenStudy (anonymous):

7sec theta - 5 sec theta = -4

OpenStudy (anonymous):

Are we solving for theta?

OpenStudy (anonymous):

solve the following trigonometric equation pu

OpenStudy (anonymous):

6 sin^2 + 3 sin theta - 3 = 0,thanks po

OpenStudy (anonymous):

pzt.thanks

OpenStudy (anonymous):

Okay. I'm assuming that you're looking for theta. --- Original Problem \[7\sec \theta-5\sec \theta=-4\] Subtract 7secant from 5secant, because they are like quantities. \[2\sec \theta=-4\] Divide both sides by two (Division Property of Equality) \[\sec \theta = -2\] Now, because secant is the inverse of cosine (or one DIVIDED by cosine) \[1/\cos \theta = -2\] We inverse both sides. \[\cos \theta=-1/2\]

OpenStudy (anonymous):

THANKS PU!!!!

OpenStudy (anonymous):

Then, we inverse cosine to find the value of theta for -1/2 \[\cos^{-1} (-.5)=1\] So, the, we can say that:\[\theta=1\] Now, depending on your question, you may have to find both answers, or all answers, but this should help you.

OpenStudy (anonymous):

merun pa pung isa 6sin^2 theta + 3 sin theta - 3 =0